Circular No. 8006

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002hi W. Li, University of California at Berkeley (UCB), reports the discov-

ery by LOTOSS (cf. IAUC 7906) of an apparent supernova (mag ∼ 17.9) on unfiltered KAIT images taken on Oct. 21.5 and 31.5 UT. SN 2002hi is located at α = 7^{h}19^{m}54^{s}.08, δ = +17^{o}58′18′′.2 (equinox 2000.0), which is within 1′′.0 of the nucleus of a diffuse faint galaxy. A KAIT image taken on Mar. 30.2 UT showed nothing at this position (limiting mag ∼ 19.0).

R. Chornock, S. Jha, and A. V. Filippenko, UCB; and B. Barris, University of Hawaii, write that a spectrum (range 400–1000 nm), obtained on Nov. 1.6 UT with the Keck II 10-m telescope (+ ESI), shows that SN 2002hi is a type-IIn supernova, with strong Hα emission having very broad (FWHM ∼ 13000 km/s) and intermediate-width (FWHM ∼ 2000 km/s) components. The very broad component is also visible in the He I 587.6-nm emission line. Intermediate-width emission similar to that of Wolf-Rayet stars is present at a rest wavelength of ∼ 465 nm, unusual even among type-IIn supernovae. Aside from the Wolf-Rayet lines, the spectrum is similar to that of SN 1988Z at early times (Filippenko 1997, ARAA 35, 309, and references therein). Accordingly, the object might become quite luminous at radio and x-ray wavelengths, because of interaction of the ejecta with dense circumstellar gas. The host-galaxy redshift derived from narrow [N II] 658.3-nm emission is 0.061.

SUPERNOVA 2002hj Further to IAUC 7999, W. M. Wood-Vasey, G. Aldering, P. Nugent,

and K. Li report the discovery of an apparent supernova (mag 17.4) on unfiltered NEAT images taken with the Haleakala 1.2-m telescope on Oct. 20.48, 20.49, and 20.50 UT. Confirming images were taken with the Palomar 1.2-m telescope on Oct. 29.40, 29.43, and 29.46 (mag 17.4). The new object is located at α = 2^{h}58^{m}09^{s}.34, δ = +4^{o}41′03′′.8 (equinox 2000.0) which is 2′′.0 west and 6′′.5 south of the center of the apparent host galaxy. No point source is visible at this location in NEAT Palomar images taken on Oct. 9 (limiting mag 19.5 at S/N = 10).

V4744 SAGITTARII Precise position measured by H. Yamaoka, Kyushu University, from

images taken by S. Kiyota and by P. Nelson: α = 17^{h}47^{m}21^{s}.724 ± 0^{s}.003, δ = −23^{o}28′22′′.79 ± 0′′.06 (equinox 2000.0; 60–70 UCAC1 reference stars).

2002 November 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green